Nelson Heating and Ventilating Company estimates the coming​ year's sales revenue based on external data. The​ company's main business is shopping mall​ construction, and it uses the square footage of each mall as a​ "yardstick" for many financial statements and projections. The company does business in four Midwest states. Last​ year, it completed heating and ventilating systems in four shopping malls with an average size of 3,090,000 square feet for sales revenues of ​$9,582,000. Nelson is the primary contractor for​ one-third of the new malls in the​ four-state area. This coming​ year, 9 new malls are under construction with an average size of 4,498,000 square feet. What is​ Nelson's anticipated sales revenue for the coming​ year?
Answer #1

Nelson will construct = (1/3) x 9 malls = 3 malls

Per square-feet price = sales revenue / total size

= $9,582,000 / (3,090,000 x 4) = $0.7752

Anticipated Sales Revenue = Average Size x No. of Malls x Per square-feet price

= 4,498,000 x 3 x $0.7752 = $10,461,125.24